# Test-Time Anchoring for Discrete Diffusion Posterior Sampling

### v1

@@ Abstract
We study the problem of posterior sampling using pretrained discrete diffusion foundation models, aiming to recover images from noisy measurements without retraining task-specific models. While diffusion models have achieved remarkable success in generative modeling, most advances rely on continuous Gaussian diffusion. In contrast, discrete diffusion offers a unified framework for jointly modeling categorical data such as text and images. Beyond unification, discrete diffusion provides faster inference, finer control, and principled training-free Bayesian inference, making it particularly well-suited for posterior sampling. However, existing approaches to discrete diffusion posterior sampling face severe challenges: derivative-free guidance yields sparse signals, continuous relaxations limit applicability, and split Gibbs samplers suffer from the curse of dimensionality. To overcome these limitations, we introduce Anchored Posterior Sampling (APS) for masked diffusion foundation models, built on two key innovations—quantized expectation for gradient-like guidance in discrete embedding space, and anchored remasking for adaptive decoding. Our approach achieves state-of-the-art performance among discrete diffusion samplers across linear and nonlinear inverse problems on the standard benchmarks. We further demonstrate the benefits of our approach in training-free stylization and text-guided editing.
@@ 1 Introduction
Diffusion models have become the state-of-the-art across a wide range of generative tasks, including images (Ramesh et al., 2021; Rombach et al., 2022; Baldridge et al., 2024; Esser et al., 2024; Black Forest Labs, 2024), audio (Huang et al., 2023; Veo, 2025), and video (Singer et al., 2023; OpenAI, 2024; Veo, 2025). Most of this progress has been driven by continuous diffusion models, where Gaussian noise is gradually added in pixel or latent space and then reversed by a learned denoiser (Sohl-Dickstein et al., 2015; Ho et al., 2020). Recently, however, discrete diffusion has emerged as a powerful alternative, showing superior performance in modeling categorical distributions such as text (Lou et al., 2024; Sahoo et al., 2024; Shi et al., 2024; Nie et al., 2025; Rout et al., 2025a) and images (Shi et al., 2024; Yang et al., 2025). Discrete diffusion further enables a unified framework for both image and text generation, supporting multimodal generation and editing.
@@
Beyond unification, discrete diffusion offers several advantages over continuous diffusion that are particularly relevant for posterior sampling. First, it achieves faster inference, often generating high-quality samples in significantly fewer reverse steps (Shi et al., 2024; Schiff et al., 2025; Ma et al., 2025). Second, it provides finer control: the model predicts a normalized categorical distribution per token (e.g., a pixel or a patch), which decouples different parts of the image, unlike Gaussian diffusion where the entire image is coupled. Third, it enables training-free posterior sampling: since the model outputs full conditional distributions at each step, these can be reweighted by the likelihood to yield a better posterior estimate posterior (Murata et al., 2024; Chu et al., 2025). This property unlocks precise image editing and solving inverse problems without additional training (§4), motivating our approach to use discrete diffusion model (Yang et al., 2025) as a prior.
State-of-the-art posterior samplers (Rout et al., 2023; 2024; Chung et al., 2024; Zhang et al., 2025) use continuous diffusion as a prior. These approaches rely on guiding the reverse diffusion process using likelihood gradients in continuous latent spaces (Rout et al., 2023; 2024; Chung et al., 2024; Zhang et al., 2025). This is infeasible for discrete diffusion due to non-differentiability in token space. Derivative-free discrete methods (Li et al., 2024) inspired by reinforcement learning provide weak signals. G2D2 (Murata et al., 2024) uses Gumbel-softmax relaxation but it is limited to discrete tokens with continuous embeddings. SGDD (Chu et al., 2025) introduces a split Gibbs sampler, but suffers from exponential complexity in sequence length (Chewi, 2023). Moreover, these methods unmask tokens in random order, which is suboptimal compared to adaptive decoding strategies in language modeling (Yang et al., 2025; Rout et al., 2025a). These limitations underscore the need for a discrete diffusion posterior sampler with adaptive decoding, leveraging next-generation multimodal models (Yang et al., 2025; Gemini Team, 2024).
Existing discrete diffusion methods explore posterior sampling under uniform (Chu et al., 2025) or mixed-noise processes (Murata et al., 2024), but these methods are tied to specific noising schemes and lack generalization (§4). Meanwhile, recent advances show that masked diffusion achieves state-of-the-art performance in image generation (Ma et al., 2025; Yang et al., 2025), yet its potential for posterior sampling remains underexplored. In this work, we take the first step towards leveraging multimodal masked diffusion models such as MMaDA (Yang et al., 2025) for posterior sampling. We introduce two key algorithmic innovations: (i) quantized expectation, which provides gradient-like guidance in purely discrete embedding space by updating the full conditional probability table (§3.2.1), and (ii) anchored remasking, an adaptive inference strategy that decodes important “anchor” tokens early in the reverse process (§3.2.2). Together, these techniques overcome the non-differentiability challenge in discrete diffusion and yield scalable posterior samplers. Extensive experiments (§4) show up to 31.36% LPIPS and 7.05% PSNR improvements on linear and nonlinear inverse problems, and illustrate training-free stylization results, as shown in Figure 1.
@@
Figure 1: We introduce Anchored Posterior Sampling (APS) for masked diffusion foundation models, built on two key innovations: (i) quantized expectation, which provides gradient-like guidance in discrete embedding space, and (ii) anchored remasking, which enables adaptive decoding. Our method supports a variety of linear and nonlinear image restoration tasks (left three columns), as well as mask-based garment styling and reference-guided style transfer (last column).
@@
  • Quantized expectation: a novel strategy to update all entries of the conditional probability table, enabling strong gradient-like guidance in the discrete embedding space (§3.2.1).
  • Anchored remasking: an adaptive decoding strategy that unmasks “anchor” tokens early in the reverse process, better utilizing model’s capacity to decode remaining tokens (§3.2.2).
[…]
2 Related Works
Gaussian pixel-space diffusion methods (Chung et al., 2023; Kawar et al., 2022; Zhu et al., 2023) address noisy inverse problems by guiding the reverse process of diffusion models trained directly in pixel space. These priors are domain-specific—for instance, a model trained on ImageNet must be used for ImageNet tasks, and one trained on FFHQ for FFHQ tasks—yielding highly informative but narrow priors. Mixing domains during training dilutes this information, and at the extreme of internet-scale training, the prior remains valid for generation (Ramesh et al., 2021; Baldridge et al., 2024; Esser et al., 2024; Black Forest Labs, 2024) but becomes less informative for specific domains. This motivates the challenge of extracting domain-relevant information from general-purpose priors.
Gaussian latent diffusion. PSLD (Rout et al., 2023) introduced posterior sampling with latent diffusion, showing how domain-specific priors can be extracted from large-scale foundation models. This initiated a new line of work (Chung et al., 2024; Song et al., 2024; Rout et al., 2024; 2025b; Noroozi et al., 2024; Zhang et al., 2025; Chung et al., 2025) leveraging the advantages of latent diffusion: a single pretrained model can handle multiple domains, enabling inverse problems and semantic edits without retraining, while also being faster and more scalable to high-resolution synthesis. A drawback, however, is that posterior sampling often requires backpropagation through large denoisers (e.g., Flux (Black Forest Labs, 2024), SD3.5 (Esser et al., 2024)), which is prohibitively slow. RB-Modulation (Rout et al., 2025c) alleviates this by framing the problem as stochastic optimal control, directly optimizing the terminal latent state and reducing runtime from several minutes (PSLD: $\sim$12 min, P2L: $\sim$30 min, STSL: $\sim$3 min) to under 40 seconds. This efficiency relies on continuous, differentiable latent embeddings, a property that does not extend to discrete diffusion. Addressing this gap motivates the need for new posterior sampling approaches in discrete settings.
Uniform discrete diffusion. Recent works have explored posterior sampling with discrete diffusion. G2D2 (Murata et al., 2024) extends the proximal sampler of RB-Modulation to VQ-diffusion (Gu et al., 2022) using a star-shaped noising process and Gumbel-Softmax dequantization (Gumbel, 1954; Jang et al., 2017; Maddison et al., 2017). While it enables gradient guidance, G2D2 depends on continuous relaxations, requires storing log-probabilities from previous step, and struggles to generalize to purely discrete token embeddings (§4). SGDD (Chu et al., 2025) instead proposes a split Gibbs sampler with Hamming-distance reweighting and rejection sampling via Metropolis–Hastings, but its exponential rejection rate restricts results to low-resolution tasks.
Masked (absorbing) discrete diffusion. While G2D2 and SGDD can, in principle, be adapted to masked diffusion, they perform poorly with purely discrete token embeddings. In contrast, our method leverages a unified masked discrete diffusion model and introduces two key components: quantized expectation (§3.2.1) and anchored remasking (§3.2.2). Together, these yield an efficient and scalable posterior sampler for high-resolution inverse problems. To our knowledge, this is the first inverse problem solver tailored for masked discrete diffusion with purely discrete embeddings, outperforming prior discrete samplers and remaining competitive—often superior—to continuous diffusion methods at substantially lower inference cost (§4).
3 Method
3.1 Preliminaries
@@
Codebook entries may be continuous vectors ${\mathbf{c}}_{j}\in\mathbb{R}^{d}$ or purely discrete binary embeddings ${\mathbf{c}}_{j}\in\{-1,+1\}^{d}$ used in this paper. Binary embeddings are particularly appealing because prior studies (Yu et al., 2024) show that masked diffusion models degrade in generation quality as continuous vocabulary size grows, yet lookup-free quantization (LFQ) with binary embeddings achieves both strong generation and reconstruction quality. In LFQ, the codebook is not learned but obtained by thresholding the encoder output:
${\mathcal{Q}}_{\mathrm{lfq}}({\mathbf{e}}^{l})\coloneqq\text{sign}({\mathbf{e}}^{l}),~\text{where}~[{\mathcal{Q}}_{\mathrm{lfq}}({\mathbf{e}}^{l})]_{i}=\begin{cases}+1&\text{if }{\mathbf{e}}^{l}[i]>0,\\ -1&\text{otherwise}.\end{cases}$ (2)
@@
Notation: We use $\coloneqq$’ to indicate architectural and parameterization choices in our model, to distinguish from ‘==’ which are identities that follow from mathematical derivations.
@@
Masked Diffusion. We now describe masked discrete diffusion, which defines a generative model over the discrete state space ${\mathcal{S}}={\mathcal{V}}^{L}$, where ${\mathcal{V}}=\{1,\dots,K,K+1\}$ consisting of $K$ codebook indices and a special [MASK] token $\mathbf{m}$} corresponding to index $K+1$. Let $X=\{x^{l}\}_{l=1}^{L}\in{\mathcal{S}}$ denote a sequence of tokens (or equivalently its one-hot representation ${\mathbf{x}}=\{{\mathbf{x}}^{l}\}_{l=1}^{L}$). The data distribution is denoted by $q(\cdot)$ over $\mathcal{S}$}. The goal of masked diffusion is to learn a generative model that samples from $q(\cdot)$. Masked diffusion models (MDMs) (Austin et al., 2021; Lou et al., 2024; Shi et al., 2024; Sahoo et al., 2024) construct a discrete-time Markov chain with $T$ steps, parameterized by $\alpha_{t}$ with $t\in[0,1]$. The forward process gradually replaces each token with the [MASK]token:
$q({\mathbf{z}}_{t}|{\mathbf{x}})=\prod_{l=1}^{L}q({\mathbf{z}}_{t}^{l}|{\mathbf{x}}),\quad q({\mathbf{z}}_{t}^{l}|{\mathbf{x}})=\mathrm{Cat}\!\left({\mathbf{z}}_{t}^{l};\alpha_{t}{\mathbf{x}}^{l}+(1-\alpha_{t}){\mathbf{m}}\right),$ (3)
@@
where ${\mathbf{z}}_{t}^{l}$ is either preserved from ${\mathbf{x}}^{l}$ with probability $\alpha_{t}$ or replaced with $\mathbf{m}$} otherwise. The corresponding reverse process is parameterized by a neural network $p_{\theta}$, which predicts categorical distributions over tokens. For each token position $l$, the transition distribution is
@@
$p_{\theta}({\mathbf{z}}_{s}^{l}|{\mathbf{z}}_{t})\coloneqq q({\mathbf{z}}_{s}^{l}|{\mathbf{z}}_{t}^{l},{\mathbf{x}}_{\theta}({\mathbf{z}}_{t}))=\begin{cases}\mathrm{Cat}({\mathbf{z}}_{s}^{l};{\mathbf{z}}_{t}^{l}),&{\mathbf{z}}_{t}^{l}\neq{\mathbf{m}},\\ \mathrm{Cat}\!\left({\mathbf{z}}_{s}^{l};\frac{\alpha_{s}-\alpha_{t}}{1-\alpha_{t}}{\mathbf{x}}_{\theta}({\mathbf{z}}_{t})+\frac{1-\alpha_{s}}{1-\alpha_{t}}{\mathbf{m}}\right),&{\mathbf{z}}_{t}^{l}={\mathbf{m}},\end{cases}$ (4)
@@
where ${\mathbf{x}}_{\theta}({\mathbf{z}}_{t})$ denotes the network prediction. Training minimizes the negative evidence lower bound (NELBO) by aligning the reverse transition $p_{\theta}({\mathbf{z}}_{s}^{l}|{\mathbf{z}}_{t})$ with the inference posterior $q({\mathbf{z}}_{s}^{l}|{\mathbf{z}}_{t}^{l},{\mathbf{x}})$ derived from the forward process (3). Concretely, the training objective ${\mathcal{L}}_{\mathrm{NELBO}}({\mathbf{x}};\theta)\coloneqq$
@@
$\mathbb{E}_{Z_{0}\sim q(\cdot|{\mathbf{x}})}\big[-\log p_{\theta}({\mathbf{x}}|Z_{0})\big]+\sum_{i=1}^{T}\mathbb{E}_{Z_{t}\sim q(\cdot|{\mathbf{x}})}\Big[\frac{\alpha_{t}-\alpha_{s}}{1-\alpha_{t}}\sum_{l=1}^{L}\log\langle{\mathbf{x}}^{l}_{\theta}(Z_{t}),{\mathbf{x}}^{l}\rangle\bm{1}_{\{Z_{t}^{l}={\mathbf{m}}\}}\Big],$ (5)
where, for brevity, we drop $i$ from $t(i)=i/T$ and s(i)=(i1)/Ts(i)=(i-1)/T.
3.2 Test-Time Anchored Posterior Sampling
This section develops our theoretical framework for posterior sampling with discrete diffusion. We first derive a training-based objective (Theorem 3.1) that incorporates measurements into the reverse diffusion process, and then present a lightweight test-time training bound (Theorem 3.2) that reuses a pretrained denoiser without retraining. Building on this, we introduce two key mechanisms—Quantized Expectation3.2.1) for differentiable likelihood evaluation and Anchored Remasking3.2.2) for adaptive unmasking—that enable scalable and accurate posterior sampling.

### v2

@@ Abstract
While continuous diffusion models have achieved remarkable success, discrete diffusion offers a unified framework for jointly modeling text and images. Beyond unification, discrete diffusion provides faster inference, finer control, and principled training-free guidance, making it well-suited for posterior sampling. Existing approaches to posterior sampling using discrete diffusion face severe challenges: derivative-free guidance yields sparse signals, continuous relaxations limit applicability, and split Gibbs samplers suffer from the curse of dimensionality. To overcome these limitations, we introduce Anchored Posterior Sampling (APS), built on two key innovations: quantized expectation for gradient-like guidance in discrete embedding space, and anchored remasking for adaptive decoding. APS achieves state-of-the-art performance among discrete diffusion samplers on both linear and nonlinear inverse problems across the standard image benchmarks. We demonstrate the generality of APS through training-free stylization and text-guided editing. We further apply APS to a large-scale diffusion language model, showing consistent improvement in question answering.
Machine Learning, ICML
[…]
@@ 1 Introduction
Diffusion models have become the state-of-the-art across a wide range of generative tasks, including images (Ramesh et al., 2021; Rombach et al., 2022; Baldridge et al., 2024; Esser et al., 2024; Black Forest Labs, 2024), audio (Huang et al., 2023; Veo, 2025), and video (Singer et al., 2023; OpenAI, 2024; Veo, 2025). Most of this progress has been driven by continuous diffusion models, where Gaussian noise is gradually added in pixel or latent space and then reversed by a learned denoiser (Sohl-Dickstein et al., 2015; Ho et al., 2020). Recently, however, discrete diffusion has emerged as a powerful alternative, showing superior performance in modeling categorical distributions such as text (Lou et al., 2024; Sahoo et al., 2024; Shi et al., 2024; Nie et al., 2025b; Rout et al., 2025a) and images (Shi et al., 2024; Yang et al., 2025b). Discrete diffusion further enables a unified framework for both image and text generation, supporting multimodal generation and editing.
@@
Beyond unification, discrete diffusion offers several advantages over continuous diffusion that are particularly relevant for posterior sampling. First, it achieves faster inference, often generating high-quality samples in significantly fewer reverse steps (Shi et al., 2024; Schiff et al., 2025; Ma et al., 2025). Second, it provides finer control: the model predicts a normalized categorical distribution per token (e.g., a pixel or a patch), which decouples different parts of the image, unlike Gaussian diffusion where the entire image is coupled. Third, it enables training-free posterior sampling: since the model outputs full conditional distributions at each step, these can be reweighted by the likelihood to yield a better posterior estimate (Murata et al., 2024; Chu et al., 2025). This property unlocks precise image editing and inverse problem solving without additional training (§4), motivating the use of discrete diffusion as a prior.
@@
Figure 1: We introduce Anchored Posterior Sampling (APS), built on two key innovations: (i) quantized expectation, which provides gradient-like guidance in discrete embedding space, and (ii) anchored remasking, which enables adaptive decoding. APS supports a variety of linear and nonlinear restoration tasks (left three columns), as well as garment styling and style transfer (fourth column). Furthermore, APS improves question answering by guiding a 7B-parameter diffusion language model without additional training (last column).
State-of-the-art posterior samplers (Rout et al., 2024; Chung et al., 2024; Zhang et al., 2025) use continuous diffusion as a prior. These approaches rely on guiding the reverse diffusion process using likelihood gradients in continuous latent spaces (Rout et al., 2023, 2024; Chung et al., 2024; Zhang et al., 2025). This is infeasible for discrete diffusion due to non-differentiability in token space. Derivative-free discrete methods (Li et al., 2024) inspired by reinforcement learning provide weak guidance. G2D2 (Murata et al., 2024) uses Gumbel-Softmax relaxation but it is limited to discrete tokens with continuous embeddings. SGDD (Chu et al., 2025) introduces a split Gibbs sampler, but suffers from exponential complexity in sequence length (Chewi, 2023). These methods unmask tokens in random order, which is suboptimal compared to adaptive decoding strategies in language modeling (Yang et al., 2025b; Rout et al., 2025a). These limitations underscore the need for a discrete diffusion posterior sampler with adaptive decoding, leveraging next-generation multimodal models (Yang et al., 2025b; Gemini Team, 2024).
In this work, we take the first step towards leveraging multimodal discrete diffusion models (Yang et al., 2025b; Ye et al., 2025) for posterior sampling. We introduce two key algorithmic innovations: (i) quantized expectation, which provides gradient-like guidance in purely discrete embedding space by updating the full conditional probability table (§3.1), and (ii) anchored remasking, an adaptive decoding strategy that unmasks important “anchor” tokens early during inference (§3.2). Together, these techniques enable discrete diffusion posterior sampling. Like classifier-free guidance (Ho & Salimans, 2022) and Tweedie-based posterior sampling (Chung et al., 2023; Rout et al., 2023) in continuous diffusion, our approach sacrifices asymptotic correctness in favor of scalable posterior inference. As illustrated in Figure 1, we show up to 35.82% LPIPS and 10.94% PSNR improvements on linear and nonlinear inverse problems including training-free stylization (§4) . We further demonstrate generalization of our algorithm to diffusion language models achieving up to 21.99% gain.
@@
  • Quantized expectation: a novel strategy to update all entries of the conditional probability table, enabling strong gradient-like guidance in discrete diffusion (§3.1).
  • Anchored remasking: an adaptive decoding strategy to unmask “anchor” tokens early that better utilizes the model’s capacity to decode remaining tokens (§3.2).
  • Extensive evaluation on linear (super resolution, Gaussian deblurring, inpainting, motion deblur) and nonlinear (HDR, nonlinear deblurring) inverse problems on FFHQ and ImageNet. We further demonstrate training-free stylization and diffusion language model guidance, highlighting flexibility beyond typical inverse problems (§4).
2 Preliminaries
@@
Codebook entries may be continuous vectors ${\mathbf{c}}_{j}\in\mathbb{R}^{d}$ or discrete binary embeddings ${\mathbf{c}}_{j}\in\{-1,+1\}^{d}$ used in this paper. Binary embeddings are particularly appealing because prior studies (Yu et al., 2024) show that masked diffusion models degrade in generation quality as continuous vocabulary size grows, yet lookup-free quantization (LFQ) with binary embeddings achieves both strong generation and reconstruction quality. In LFQ, the codebook is not learned but obtained by thresholding the encoder output: ${\mathcal{Q}}_{\mathrm{lfq}}({\mathbf{e}}^{l})\coloneqq\text{sign}({\mathbf{e}}^{l}),$
$[{\mathcal{Q}}_{\mathrm{lfq}}({\mathbf{e}}^{l})]_{i}=\begin{cases}+1&\text{if }{\mathbf{e}}^{l}[i]>0,\\ -1&\text{otherwise}.\end{cases}$ (2)
@@
Notation: We use $\coloneqq$’ to indicate architectural and parameterization choices, to distinguish from ‘==’ which is an identity that follows from mathematical derivations.
@@
Masked diffusion defines a generative model over the discrete state space ${\mathcal{S}}={\mathcal{V}}^{L}$, where ${\mathcal{V}}=\{1,\dots,K,K+1\}$ consisting of $K$ codebook indices and a special [MASK] token $\mathbf{m}$} corresponding to index $K+1$. Let $X=\{x^{l}\}_{l=1}^{L}\in{\mathcal{S}}$ denote a sequence of tokens (or equivalently its one-hot representation ${\mathbf{x}}=\{{\mathbf{x}}^{l}\}_{l=1}^{L}$). The data distribution is denoted by $q(\cdot)$ over $\mathcal{S}$}. The goal of masked diffusion is to learn a generative model that samples from $q(\cdot)$. Masked diffusion models (MDMs) (Austin et al., 2021; Lou et al., 2024; Sahoo et al., 2024; Shi et al., 2024) construct a discrete-time Markov chain with $T$ steps, parameterized by $\alpha_{t}$ with $t\in[0,1]$. The forward process gradually replaces each token with the [MASK]token: $q({\mathbf{z}}_{t}|{\mathbf{x}})=\prod_{l=1}^{L}q({\mathbf{z}}_{t}^{l}|{\mathbf{x}}),$
$q({\mathbf{z}}_{t}^{l}|{\mathbf{x}})=\mathrm{Cat}\!\left({\mathbf{z}}_{t}^{l};\alpha_{t}{\mathbf{x}}^{l}+(1-\alpha_{t}){\mathbf{m}}\right),$ (3)
@@
where ${\mathbf{z}}_{t}^{l}$ is either preserved from ${\mathbf{x}}^{l}$ with probability $\alpha_{t}$ or replaced with $\mathbf{m}$}. The corresponding reverse process is parameterized by a neural network $p_{\theta}$, which predicts categorical distributions over tokens. For each token position $l$, the transition probability is $p_{\theta}({\mathbf{z}}_{s}^{l}|{\mathbf{z}}_{t})\coloneqq q({\mathbf{z}}_{s}^{l}|{\mathbf{z}}_{t}^{l},{\mathbf{x}}_{\theta}({\mathbf{z}}_{t}))=$
@@
$\begin{cases}\mathrm{Cat}({\mathbf{z}}_{s}^{l};{\mathbf{z}}_{t}^{l}),&{\mathbf{z}}_{t}^{l}\neq{\mathbf{m}},\\ \mathrm{Cat}\!\left({\mathbf{z}}_{s}^{l};\frac{\alpha_{s}-\alpha_{t}}{1-\alpha_{t}}{\mathbf{x}}_{\theta}({\mathbf{z}}_{t})+\frac{1-\alpha_{s}}{1-\alpha_{t}}{\mathbf{m}}\right),&{\mathbf{z}}_{t}^{l}={\mathbf{m}},\end{cases}$ (4)
@@
where ${\mathbf{x}}_{\theta}({\mathbf{z}}_{t})$ denotes the network prediction. Training minimizes the negative evidence lower bound (NELBO) by aligning the reverse transition $p_{\theta}({\mathbf{z}}_{s}^{l}|{\mathbf{z}}_{t})$ with the inference posterior $q({\mathbf{z}}_{s}^{l}|{\mathbf{z}}_{t}^{l},{\mathbf{x}})$ derived from (3). Concretely, the training objective ${\mathcal{L}}_{\mathrm{NELBO}}({\mathbf{x}};\theta)\coloneqq\mathbb{E}_{Z_{0}\sim q(\cdot|{\mathbf{x}})}\big[-\log p_{\theta}({\mathbf{x}}|Z_{0})\big]+$
@@
$\sum_{i=1}^{T}\mathbb{E}_{Z_{t}\sim q(\cdot|{\mathbf{x}})}\!\Big[\!\frac{\alpha_{t}-\alpha_{s}}{1-\alpha_{t}}\!\sum_{l=1}^{L}\!\log\langle{\mathbf{x}}^{l}_{\theta}(Z_{t}),\!{\mathbf{x}}^{l}\rangle\bm{1}_{\{Z_{t}^{l}={\mathbf{m}}\}}\Big]$ (5)
where, for brevity, we drop $i$ from $t(i)=i/T$ and s(i)=(i1)/Ts(i)=(i-1)/T. The objective (5) admits a score-based interpretation (Lou et al., 2024) and supports time-independent parameterization (Ou et al., 2025), which simplifies training and improves scalability (Nie et al., 2025a, b).
3 Test-Time Anchored Posterior Sampling